# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 4.1 Arrangements of rectangular prisms
where the second-last inequality used the hypothesis that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$ (and hence $\sum|R|\geq\delta^{\eta}$) and the final inequality used the hypothesis that $s\leq\delta^{\varepsilon}t$ and hence $s\leq\delta^{\varepsilon}$. If $\delta_{0}$ is sufficiently small depending on $\eta$, then Conclusion (A) holds with $r=1$, and we are done.
Step 2.
Suppose instead that $\mu\geq\delta^{-\varepsilon^{2}+3\eta}(st\#\mathcal{R})$. Each $s\times t\times 1$ prism $R\in\mathcal{R}$ is the image of the set $[-s/2,s/2]\times[-t/2,t/2]\times[-1/2,1/2]$ under a rigid motion $\tau$. We define $\ell=\ell(R)$ to be the image of the $z$-axis under $\tau$, and we define $\Pi(R)$ to be the image of the $yz$-plane under $\tau$.
For each $x\in\bigcup_{R\in\mathcal{R}}Y_{2}(R)$, define the “plane spread”
$\operatorname{P-spread}(x)=\max\Big{\{}\angle\big{(}\Pi(R),\Pi(R^{\prime})\big{)}\Big{\}},$
@@
where the maximum is taken over all pairs $R,R^{\prime}$ for which $x\in Y_{2}(R)$ and $x\in Y_{2}(R^{\prime})$, and where $\angle(\Pi,\Pi^{\prime})$ denotes the unsigned angle between the normal vectors of $\Pi$ and $\Pi^{\prime}$. Since $\Pi(R)$ and $\Pi(R^{\prime})$ are defined up to uncertainty $O(s/t)$, $\operatorname{P-spread}(x)$ is meaningfully defined if it is substantially larger than $s/t$. If $x\not\in\bigcup_{R\in\mathcal{R}}Y_{2}(R)$, define $\operatorname{P-spread}(x)=0$.
@@
Step 3. Suppose there exists a prism $R_{0}\in\mathcal{R}$ and a number $\theta\geq\delta^{-4\eta/\varepsilon}\frac{s}{t}$ so that
@@
|$\big{|}\{x\in R_{0}\colon\theta\leq\operatorname{P-spread}(x)<2\theta\}\big{|}\geq\frac{1}{100}|\log\delta|^{-2}\delta^{\eta}|R_{0}|.$ (4.6)
@@
Since $R_{0}$ is a $s\times t\times 1$ prism, for each $x_{0}\in R_{0}$ we have $|R_{0}\cap B(x_{0},t)|\sim st^{2}$. By pigeonholing, we can select a point $x_{0}\in R_{0}$ so that
@@
|$\big{|}\{x\in R_{0}\cap B(x_{0},t)\colon\theta\leq\operatorname{P-spread}(x)<2\theta\}\big{|}\gtrapprox\delta^{\eta}st^{2}.$ (4.7)
@@
After a harmless translation and rotation, we may suppose that $\Pi(R_{0})$ is the $yz$-plane, i.e. $\operatorname{span}\{{\bf e}_{2},{\bf e}_{3}\}$, and that $R_{0}\cap B(x_{0},t)$ is contained in the “square” $S=[-s,s]\times[-t,t]\times[-t,t]$ (note as well that $|S|\sim|R_{0}\cap B(x_{0},t)|$). Let $\tilde{S}$ denote the thickened square $[-10t\theta,10t\theta]\times[-10t,10t]\times[-10t,10t]$, i.e. $\tilde{S}$ is comparable to the $t\theta$ neighborhood of $S$.
@@
Denote the set on the LHS of (4.7) by $E$; then for each $x\in E$ there is a prism $R_{x}\in\mathcal{R}$ with $x\in Y_{2}(R)$ and $\angle(\Pi(R_{x}),\operatorname{span}\{{\bf e}_{2},{\bf e}_{3}\})\in[\theta,2\theta)$. This implies that $R_{x}\cap B(x,t)\subset R_{x}\cap B(x_{0},10t)\subset\tilde{S}$. Since the shading $Y_{2}(R_{x})$ is regular, we also have
$|Y_{2}(R_{x})\cap B(x_{0},2t)|\gtrapprox|Y_{2}(R_{x})|\frac{|R_{x}\cap B(x_{0},2t)|}{|R_{x}|}\gtrsim\delta^{\eta}st^{2}.$ (4.8)
Define the linear map $A(x,y,z)=\Big{(}\frac{x}{10t\theta},\ \frac{y}{10t},\ \frac{z}{10t}\Big{)}$, so $A(\tilde{S})=[-1,1]^{3}$. Define
$\hat{S}=A(S)=[-\frac{s}{10t\theta},\frac{s}{10t\theta}]\times[-1/10,1/10]\times[-1/10,1/10],$
@@
and let $\hat{E}=A(E)\subset\hat{S}$. For each $x\in E$, let $\hat{R}_{x}=A(R_{x}\cap\tilde{S})$; this is comparable to a rectangular prism of dimensions roughly $\frac{s}{t\theta}\times 1\times 1$, and this prism makes angle $\sim 1$ with the $yz$-plane, i.e. the prism has a normal direction ${\bf n}(\hat{R}_{x})$ (defined up to uncertainty $O(\frac{s}{t\theta})$) with $|{\bf n}(\hat{R}_{x})\wedge{\bf e}_{1}|\gtrsim 1$. By (4.8) we have
$|\hat{Y}_{2}(\hat{R}_{x})|\gtrapprox\delta^{\eta}\frac{s}{t\theta}.$ (4.9)
@@
By pigeonholing, we can find a unit vector $e\in\mathbb{R}^{3}$ and a set $\hat{E}^{\prime}\subset\hat{E}$ with $|\hat{E}^{\prime}|\sim|\hat{E}|$, so that $\angle\big{(}e,{\bf n}(\hat{R}_{x})\big{)}\leq\frac{1}{100}$ for each $x\in\hat{E}^{\prime}$. In particular, this implies $|e\wedge{\bf e}_{1}|\gtrsim 1$, and hence the projection of $e$ to the $yz$-plane has magnitude $\sim 1$; denote the image of $e$ under this projection by $e^{\prime}$.
@@
(4.7) says that $|E|\gtrapprox\delta^{\eta}st^{2}$ and hence $|\hat{E}^{\prime}|\gtrsim|\hat{E}|\gtrapprox\delta^{\eta}\frac{s}{t\theta}$. On the other hand, $\hat{E}^{\prime}\subset\hat{S}$, and the latter set has volume $\sim\frac{s}{t\theta}$. Since $\hat{S}=[-\frac{s}{10t\theta},\frac{s}{10t\theta}]\times[-1/10,1/10]\times[-1/10,1/10]$ and $e^{\prime}$ is contained in the $yz$ plane, by Fubini we can select a line $\ell$ pointing in direction $e^{\prime}$ that satisfies $|\ell\cap\hat{E}^{\prime}|\gtrapprox\delta^{\eta}$.
@@
Select a $\frac{s}{t\theta}$-separated set of points $x_{1},\ldots,x_{N}$, with $N\gtrapprox\delta^{\eta}\frac{t\theta}{s}$, from $|\ell\cap\hat{E}^{\prime}|$; we will index these points so that $|x_{i}-x_{j}|\geq\frac{s}{t\theta}|i-j|$. Note that $R_{x_{i}}\cap\ell$ is a line segment of length between $\frac{s}{t\theta}$ and $10\frac{s}{t\theta}$, and hence if $i,j$ are a pair of indices with $|x_{i}-x_{j}|\geq 100\frac{s}{t\theta}$, then $\operatorname{dist}(R_{x_{i}}\cap\ell,R_{x_{j}}\cap\ell)\gtrsim\frac{s}{t\theta}|i-j|$. Since $\ell$ points in direction $e^{\prime}$, and $|e^{\prime}\cdot{\bf n}(\hat{R}_{x_{j}})|\sim 1$, we conclude that $\operatorname{dist}(x_{i},R_{x_{j}})\gtrsim\frac{s}{t\theta}|i-j|$, so in particular $R_{x_{i}}$ contains a point that has distance $\gtrsim\frac{s}{t\theta}|i-j|$ from $R_{x_{j}}$. Thus either $R_{x_{i}}$ and $R_{x_{j}}$ do not intersect, or if they do intersect then their normal vectors satisfy $\angle\big{(}{\bf n}(\hat{R}_{x_{i}}),\ {\bf n}(\hat{R}_{x_{j}})\big{)}\gtrsim\frac{s}{t\theta}|i-j|.$ In either case, we have
$|\hat{R}_{x_{i}}\cap\hat{R}_{x_{j}}|\lesssim\frac{(\frac{s}{t\theta})^{2}}{\frac{s}{t\theta}(1+|i-j|)}=\frac{s}{t\theta(1+|i-j|)}.$ (4.10)
@@
The above inequality vacuously holds if $i,j$ are a pair of indices with $|x_{i}-x_{j}|<100\frac{s}{t\theta}$, since in this case $|i-j|\sim 1$ and we can use the trivial bound $|\hat{R}_{x_{i}}\cap\hat{R}_{x_{j}}|\leq|\hat{R}_{x_{i}}|$.
@@
$\Big{\|}\sum_{i=1}^{N}\chi_{\hat{Y}_{2}(\hat{R}_{i})}\Big{\|}_{2}^{2}\leq\sum_{i,j=1}^{N}|\hat{R}_{i}\cap\hat{R}_{j}|\lesssim N\frac{s}{t\theta}\big{|}\log\big{(}\frac{s}{t\theta}\big{)}\big{|},$ (4.11)
@@
$\Big{|}\bigcup_{i=1}^{N}\hat{Y}_{2}(\hat{R}_{i})\Big{|}\geq\frac{\Big{(}\sum_{i}|\hat{Y}_{2}(\hat{R}_{i})|\Big{)}^{2}}{\Big{\|}\sum_{i}\chi_{\hat{Y}_{2}(\hat{R}_{i})}\Big{\|}_{2}^{2}}\gtrapprox\frac{\Big{(}N\delta^{\eta}\frac{s}{t\theta}\Big{)}^{2}}{N\frac{s}{t\theta}\big{|}\log\big{(}\frac{s}{t\theta}\big{)}\big{|}}\gtrapprox\delta^{2\eta}N\frac{s}{t\theta}\gtrapprox\delta^{3\eta},$
@@
$\Big{|}\tilde{S}\cap\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\gtrapprox\delta^{3\eta}|\tilde{S}|.$

### v2

@@ 4.1 Arrangements of rectangular prisms
where the final inequality used the hypothesis that $s\leq\delta^{\varepsilon}t$ and hence $s\leq\delta^{\varepsilon}$. Thus Conclusion (A) holds with $r=1$ and $B$ the unit ball, and we are done.
Our choice of $\varepsilon_{1}$ will satisfy $\varepsilon_{1}\leq\varepsilon^{3}/6$, and thus for the remainder of the proof we may suppose that
$\mu>\delta^{-(2/3)\varepsilon^{2}}(st\#\mathcal{R}).$ (4.4)
Step 2: Localization. Let $\mathbb{T}_{u}$ be a set of essentially distinct $u$ tubes with the property that for every $u$-tube $T_{u}$, there is at least one, and at most $O(1),$ tubes from $\mathbb{T}_{u}$ whose 2-fold dilate contains $T_{u}$. For each $x\in\bigcup_{R\in\mathcal{R}}Y_{3}(R)$, select a tube $T_{u}(x)\in\mathbb{T}_{u}$ with the property that $T_{u_{x}}\subset 2T_{u}$. Define the shading $Y(T_{u})=\{x\in\mathbb{R}^{3}\colon T_{u}(x)=T_{u}\}$. We have $\bigcup_{R\in\mathcal{R}}Y_{3}(R)=\bigsqcup_{T_{u}\in\mathbb{T}_{u}}Y(T_{u})$. Note that technically, $Y(T_{u})$ is a subset of $2T_{u}$, rather than a subset of $T_{u}$. Observe that if $R\in\mathcal{R}$ with $x\in Y_{3}(R)$, then $R\subset T_{u_{x}}\subset 2T_{u}(x)$.
Since the tubes in $\mathbb{T}_{u}$ are essentially distinct, for each $T_{u}\in\mathbb{T}_{u}$ there are $O(1)$ tubes $T_{u}^{\prime}\in\mathbb{T}_{u}$ with the property that $2T_{u}\cap 2T_{u}^{\prime}$ contains a unit line segment — indeed, if $2T_{u}\cap 2T_{u}^{\prime}$ contains a unit line segment, then $T_{u}$ and $T_{u}^{\prime}$ must make angle $\lesssim u$. Thus by pigeonholing, we can replace $\mathbb{T}_{u}$ by a $O(1)$ refinement $\mathbb{T}_{u}^{\prime}$ with the following properties:
  • For each pair of distinct tubes $T_{u},T_{u}^{\prime}\in\mathbb{T}_{u}^{\prime}$, $2T_{u}\cap 2T_{u}^{\prime}$ does not contain a unit line segment. In particular, each $R\in\mathcal{R}$ is contained in at most one dilated tube $2T_{u}$.
  • $\Big{|}\bigsqcup_{T_{u}\in\mathbb{T}_{u}^{\prime}}Y(T_{u})\Big{|}\gtrsim\Big{|}\bigsqcup_{T_{u}\in\mathbb{T}_{u}}Y(T_{u})\Big{|}=\Big{|}\bigcup_{R\in\mathcal{R}}Y_{3}(R)\Big{|}.$
For each $R\in\mathcal{R}$, define $Y_{4}(R)=Y_{3}(R)\cap Y(T_{u})$, where $T_{u}$ is the unique tube from $\mathbb{T}_{u}^{\prime}$ for which $R\subset 2T_{u}$. If no such tube exists, then define $Y_{4}(R)=\emptyset$. We have
$\sum_{R\in\mathcal{R}}|Y_{4}(R)|\geq\mu\sum_{T_{u}\in\mathbb{T}_{u}^{\prime}}|Y(T_{u})|\gtrsim\mu\sum_{T_{u}\in\mathbb{T}_{u}}|Y(T_{u})|\geq\frac{1}{2}\sum_{R\in\mathcal{R}}|Y_{3}(R)|\geq\delta^{2\varepsilon_{1}}\sum_{R\in\mathcal{R}}|R|.$
Finally, let $\mathbb{T}_{u}^{\prime\prime}\subset\mathbb{T}_{u}^{\prime}$ consist of those $T_{u}\in\mathbb{T}_{u}^{\prime}$ for which
$\sum_{R\in\mathcal{R}[2T_{u}]}|Y_{4}(R)|\geq c\delta^{2\varepsilon_{1}}\sum_{R\in\mathcal{R}[2T_{u}]}|R|.$ (4.5)
We will choose the constant $c\sim 1$ sufficiently small so that
$\sum_{T_{u}\in\mathbb{T}_{u}^{\prime\prime}}\sum_{R\in\mathcal{R}[2T_{u}]}|Y_{4}(R)|\gtrsim\sum_{T_{u}\in\mathbb{T}_{u}^{\prime}}\sum_{R\in\mathcal{R}[2T_{u}]}|Y_{4}(R)|.$
In summary, the situation is as follows: For each $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$, we have
$\sum_{R\in\mathcal{R}[2T_{u}]}|Y_{4}(R)|\gtrsim\delta^{2\varepsilon_{1}}\sum_{R\in\mathcal{R}[2T_{u}]}|R|,$ (4.6)
and for each $x\in\bigcup_{R\in\mathcal{R}[2T_{u}]}Y_{4}(R),$ we have
$\{R\in\mathcal{R}[2T_{u}]\colon x\in Y_{4}(R)\}=\{R\in\mathcal{R}\colon x\in Y_{3}(R)\}=\{R\in\mathcal{R}\colon x\in Y_{2}(R)\}.$ (4.7)
The cardinality of the above set is between $\mu$ and $2\mu$.
For each $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$ and each $R\in\mathcal{R}[2T_{u}]$, the unit rescaling of $R$ relative to $2T_{u}$ is comparable to a $s/u\times t/u\times 1$ prism. Define $\tilde{s}=s/u$ and $\tilde{t}=t/u$. Abusing notation slightly, we will call this set a $\tilde{s}\times\tilde{t}\times 1$ prism, and we will denote the (multi) set of these prisms by ~$\tilde{\mathcal{R}}^{T_{u}}$. Let $\tilde{Y}_{1}$ (resp. $\tilde{Y}_{4}$) be the images of the shadings $Y_{1}$ (resp. $Y_{4}$) under the unit rescaling described above. Thus for each $T_{u}\in\mathbb{T}_{u}$, we have that (~$(\tilde{\mathcal{R}}^{T_{u}},\tilde{Y}_{1})$ is a set of $\tilde{s}\times\tilde{t}\times 1$ prisms and a shading on these prisms, and similarly for (~$(\tilde{\mathcal{R}}^{T_{u}},\tilde{Y}_{4})$. By (4.7) and our definition of $Y_{3}$, for each $T_{u}\in\mathbb{T}_{u}$ and each $x\in\bigcup_{\tilde{R}\in\tilde{\mathcal{R}}}\tilde{Y}_{4}(\tilde{R})$, we have
$\mu\leq\#\{\tilde{R}\in\tilde{\mathcal{R}}^{T_{u}}\colon x\in\tilde{Y}_{4}(\tilde{R})\}<2\mu.$ (4.8)
For each $\tilde{R}\in\tilde{\mathcal{R}}^{T_{u}}$, let $v(\tilde{R})$ be the direction of a tube of minimal diameter that contains $\tilde{R}$. Thus $v(R)$ is the “direction” of the tube $\tilde{R}$. We have that $v(R)$ is meaningfully defined up to accuracy $\tilde{t}$, in the sense that if $\tilde{R}$ and $\tilde{R}^{\prime}$ are comparable, then $\angle(v(\tilde{R}),\ v(\tilde{R}^{\prime}))=O(\tilde{t})$.
As a consequence of (4.3), we have that for each $x\in\bigcup_{\tilde{R}\in\tilde{\mathcal{R}}^{T_{u}}}\tilde{Y}_{4}(\tilde{R})$, each $\tilde{t}\leq r\leq 1$, and each unit vector $v$, we have
#$\#\{\tilde{R}\in\tilde{\mathcal{R}}^{T_{u}}\colon x\in\tilde{Y}_{4}(\tilde{R}),\ \angle(v(\tilde{R}),\ v)\leq r\}\lesssim r^{\varepsilon_{1}}\mu.$ (4.9)
Step 3: Dealing with transverse intersection. Fix a $u$-tube $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$, and define ~=~$\tilde{\mathcal{R}}=\tilde{\mathcal{R}}^{T_{u}}$. Each $\tilde{R}\in\tilde{\mathcal{R}}$ is comparable to a rectangular prism of dimensions $\tilde{s}\times\tilde{t}\times 1$, i.e. $\tilde{R}$ is comparable to the intersection of the following three sets: the $\tilde{s}$ neighbourhood of a plane $\Pi(\tilde{R})$; the $\tilde{t}$ neighbourhood of a tube $T(\tilde{R})$; and the unit ball. Let $n(\tilde{R})$ be the unit normal vector of $\Pi(\tilde{R})$; this is meaningfully defined up to accuracy $\tilde{s}/\tilde{t}=s/t$, in the sense that if $\tilde{R}$ and $\tilde{R}^{\prime}$ are comparable, then $\angle(n(\tilde{R}),\ n(\tilde{R}^{\prime}))=O(\tilde{s}/\tilde{t})$ (here (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ ) refers to unsigned angle, since we do not distinguish between the vectors $n(\tilde{R})$ and $-n(\tilde{R})$).
For each $x\in\bigcup_{\tilde{R}\in\tilde{\mathcal{R}}}\tilde{Y}_{4}(R)$, define the “plane spread”
$\operatorname{P-spread}(x)=\max\Big{\{}\angle\big{(}n(\tilde{R}),n(\tilde{R}^{\prime})\big{)}\Big{\}},$
@@
where the maximum is taken over all pairs $\tilde{R},\tilde{R}^{\prime}$ for which $x\in\tilde{Y}_{4}(R)$ and $x\in\tilde{Y}_{4}(R)$. Since $n(\tilde{R})$ and $n(\tilde{R}^{\prime})$ are defined up to uncertainty $O(\tilde{s}/\tilde{t})=O(s/t)$, $\operatorname{P-spread}(x)$ is meaningfully defined if it is substantially larger than $s/t$. If $x\not\in\bigcup_{\tilde{R}\in\tilde{\mathcal{R}}}\tilde{Y}_{4}(\tilde{R})$, then define $\operatorname{P-spread}(x)=0$.
@@
Suppose there exists a prism $\tilde{R}_{0}\in\tilde{\mathcal{R}}$ and a number $\theta\geq\delta^{-7\varepsilon_{1}/\varepsilon}(\tilde{s}/\tilde{t}\phantom{.})$ so that
@@
|$\big{|}\{x\in\tilde{R}_{0}\colon\theta\leq\operatorname{P-spread}(x)<2\theta\}\big{|}\geq\frac{c}{100}|\log\delta|^{-2}\delta^{2\varepsilon_{1}}|\tilde{R}_{0}|,$ (4.10)
where $c\sim 1$ is the constant from (4.5).
@@
Since $\tilde{R}_{0}$ is a $\tilde{s}\times\tilde{t}\times 1$ prism, for each $x_{0}\in\tilde{R}_{0}$ we have $|\tilde{R}_{0}\cap B(x_{0},\tilde{t})|\sim\tilde{s}\tilde{t}\phantom{.}^{2}$. By pigeonholing, we can select a point $x_{0}\in\tilde{R}_{0}$ so that
@@
|$\big{|}\{x\in\tilde{R}_{0}\cap B(x_{0},\tilde{t})\colon\theta\leq\operatorname{P-spread}(x)<2\theta\}\big{|}\gtrapprox\delta^{2\varepsilon_{1}}\tilde{s}\tilde{t}\phantom{.}^{2}.$ (4.11)
@@
After a harmless translation and rotation, we may suppose that $\Pi(\tilde{R}_{0})$ is the $yz$-plane, i.e. $\operatorname{span}\{e_{2},e_{3}\}$, and that $\tilde{R}_{0}\cap B(x_{0},\tilde{t})$ is contained in the “square” $S=[-\tilde{s},\tilde{s}]\times[-\tilde{t},\tilde{t}\phantom{.}]\times[-\tilde{t},\tilde{t}\phantom{.}]$ (note as well that $|S|\sim|\tilde{R}_{0}\cap B(x_{0},\tilde{t})|$). Let $S^{{\dagger}}$ denote the thickened square $[-10\tilde{t}\theta,10\tilde{t}\theta]\times[-10\tilde{t},10\tilde{t}\phantom{.}]\times[-10\tilde{t},10\tilde{t}\phantom{.}]$, i.e. $S^{{\dagger}}$ is comparable to the $\tilde{t}\theta$ neighborhood of $S$.
@@
Denote the set on the LHS of (4.11) by $E$; then for each $x\in E$ there is a prism $\tilde{R}_{x}\in\mathcal{R}$ with $x\in\tilde{Y}_{4}(\tilde{R})$ and $\angle(n(\tilde{R}_{x}),e_{1})\in[\theta,2\theta)$. This implies that $\tilde{R}_{x}\cap B(x,\tilde{t})\subset\tilde{R}_{x}\cap B(x_{0},10\tilde{t})\subset S^{{\dagger}}$. Recall that the shading $\tilde{Y}_{1}(\tilde{R}_{x})$ is regular, and thus we have
$|\tilde{Y}_{1}(\tilde{R}_{x})\cap B(x_{0},2\tilde{t})|\gtrapprox|\tilde{Y}_{1}(\tilde{R}_{x})|\frac{|\tilde{R}_{x}\cap B(x_{0},2\tilde{t})|}{|\tilde{R}_{x}|}\gtrsim\delta^{2\varepsilon_{1}}\tilde{s}\tilde{t}\phantom{.}^{2}.$ (4.12)
Define the linear map $\tilde{A}(x,y,z)=\Big{(}\frac{x}{10\tilde{t}\theta},\ \frac{y}{10\tilde{t}},\ \frac{z}{10\tilde{t}}\Big{)}$, so $A(S^{{\dagger}})=[-1,1]^{3}$. Define
$\hat{S}=A(S)=[-\frac{\tilde{s}}{10\tilde{t}\theta},\frac{\tilde{s}}{10\tilde{t}\theta}]\times[-1/10,1/10]\times[-1/10,1/10],$
@@
and let $\hat{E}=A(E)\subset\hat{S}$. For each $x\in E$, let $\hat{R}_{x}=A(\tilde{R}_{x}\cap S^{\dagger})$; this is comparable to a rectangular prism of dimensions roughly $\frac{\tilde{s}}{\tilde{t}\theta}\times 1\times 1$, and this prism makes angle $\sim 1$ with the $yz$-plane, i.e. there is a constant $c_{0}\sim 1$ (independent of the choice of $x\in E$) so that the prism has a normal direction $n(\hat{R}_{x})$ (defined up to uncertainty $O(\frac{\tilde{s}}{\tilde{t}\theta})$) with $\angle(n(\hat{R}_{x}),e_{1})\geq c_{0}$. By (4.12) we have
$|\hat{Y}_{2}(\hat{R}_{x})|\gtrapprox\delta^{2\varepsilon_{1}}\frac{\tilde{s}}{\tilde{t}\theta}.$ (4.13)
@@
By pigeonholing, we can find a unit vector $e\in\mathbb{R}^{3}$ and a set $\hat{E}^{\prime}\subset\hat{E}$ with $|\hat{E}^{\prime}|\sim|\hat{E}|$, so that $\angle\big{(}e,n(\hat{R}_{x})\big{)}\leq c_{0}/2$ for each $x\in\hat{E}^{\prime}$. In particular this means that $\angle(e,e_{1})\geq c_{0}/2\gtrsim 1$, and hence the projection of $e$ to the $yz$-plane has magnitude $\sim 1$; denote the image of $e$ under this projection by $e^{\prime}$.
@@
(4.11) says that $|E|\gtrapprox\delta^{2\varepsilon_{1}}\tilde{s}\tilde{t}\phantom{.}^{2}$ and hence $|\hat{E}^{\prime}|\gtrsim|\hat{E}|\gtrapprox\delta^{\eta}\frac{s}{t\theta}$. On the other hand, $\hat{E}^{\prime}\subset\hat{S}$, and the latter set has volume $\sim\frac{\tilde{s}}{\tilde{t}\theta}$. Since $\hat{S}=[-\frac{\tilde{s}}{10\tilde{t}\theta},\frac{\tilde{s}}{10\tilde{t}\theta}]\times[-1/10,1/10]\times[-1/10,1/10]$ and $e^{\prime}$ is contained in the $yz$ plane, by Fubini we can select a line $\ell$ pointing in direction $e^{\prime}$ that satisfies $|\ell\cap\hat{E}^{\prime}|\gtrapprox\delta^{2\varepsilon_{1}}$.
@@
Select a $\frac{\tilde{s}}{\tilde{t}\theta}$-separated set of points $x_{1},\ldots,x_{N}$ (here $N\gtrapprox\delta^{2\varepsilon_{1}}\frac{\tilde{t}\theta}{\tilde{s}}$) from $|\ell\cap\hat{E}^{\prime}|$. We will index these points so that $|x_{i}-x_{j}|\geq\frac{\tilde{s}}{\tilde{t}\theta}|i-j|$. Note that $\hat{R}_{x_{i}}\cap\ell$ is a line segment of length between $\frac{\tilde{s}}{\tilde{t}\theta}$ and $10\frac{\tilde{s}}{\tilde{t}\theta}$, and hence if $i,j$ is a pair of indices with $|x_{i}-x_{j}|\geq 100\frac{\tilde{s}}{\tilde{t}\theta}$, then $\operatorname{dist}(\hat{R}_{x_{i}}\cap\ell,\hat{R}_{x_{j}}\cap\ell)\gtrsim\frac{\tilde{s}}{\tilde{t}\theta}|i-j|$. Since the line $\ell$ points in direction $e^{\prime}$, and $|e^{\prime}\cdot n(\hat{R}_{x_{j}})|\sim 1$, we conclude that $\operatorname{dist}(x_{i},\hat{R}_{x_{j}})\gtrsim\frac{\tilde{s}}{\tilde{t}\theta}|i-j|$, so in particular $\hat{R}_{x_{i}}$ contains a point that has distance $\gtrsim\frac{\tilde{s}}{\tilde{t}\theta}|i-j|$ from $\hat{R}_{x_{j}}$. Thus either $\hat{R}_{x_{i}}$ and $\hat{R}_{x_{j}}$ do not intersect, or if they do intersect then their normal vectors satisfy $\angle\big{(}n(\hat{R}_{x_{i}}),\ n(\hat{R}_{x_{j}})\big{)}\gtrsim\frac{\tilde{s}}{\tilde{t}\theta}|i-j|.$ In either case, we have
$|\hat{R}_{x_{i}}\cap\hat{R}_{x_{j}}|\lesssim\frac{(\frac{\tilde{s}}{\tilde{t}\theta})^{2}}{\frac{\tilde{s}}{\tilde{t}\theta}(1+|i-j|)}=\frac{\tilde{s}}{\tilde{t}\theta(1+|i-j|)}.$ (4.14)
@@
The above inequality vacuously holds if $i,j$ are a pair of indices with $|x_{i}-x_{j}|<100\frac{\tilde{s}}{\tilde{t}\theta}$, since in this case $|i-j|\sim 1$ and we can use the trivial bound $|\hat{R}_{x_{i}}\cap\hat{R}_{x_{j}}|\leq|\hat{R}_{x_{i}}|$.
@@
$\Big{\|}\sum_{i=1}^{N}\chi_{\hat{Y}_{2}(\hat{R}_{i})}\Big{\|}_{2}^{2}\leq\sum_{i,j=1}^{N}|\hat{R}_{i}\cap\hat{R}_{j}|\lesssim N\frac{\tilde{s}}{\tilde{t}\theta}\big{|}\log\big{(}\frac{\tilde{s}}{\tilde{t}\theta}\big{)}\big{|},$ (4.15)
@@
$\Big{|}\bigcup_{i=1}^{N}\hat{Y}_{2}(\hat{R}_{i})\Big{|}\geq\frac{\Big{(}\sum_{i}|\hat{Y}_{2}(\hat{R}_{i})|\Big{)}^{2}}{\Big{\|}\sum_{i}\chi_{\hat{Y}_{2}(\hat{R}_{i})}\Big{\|}_{2}^{2}}\gtrapprox\frac{\Big{(}N\delta^{2\varepsilon_{1}}\frac{\tilde{s}}{\tilde{t}\theta}\Big{)}^{2}}{N\frac{\tilde{s}}{\tilde{t}\theta}\big{|}\log\big{(}\frac{\tilde{s}}{\tilde{t}\theta}\big{)}\big{|}}\gtrapprox\delta^{4\varepsilon_{1}}N\frac{\tilde{s}}{\tilde{t}\theta}\gtrapprox\delta^{6\varepsilon_{1}},$
@@
$\Big{|}S^{\dagger}\cap\bigcup_{\tilde{R}\in\tilde{\mathcal{R}}}\tilde{Y}_{1}(\tilde{R})\Big{|}\gtrapprox\delta^{6\varepsilon_{1}}|S^{\dagger}|.$